OpenStudy (princesssleelee):

Given the expression 3x2y + 4xy + 1 – 5x4y – 6y, do the following as instructed below: Write the polynomial in descending order. Classify the polynomial by the number of terms. State the degree of the polynomial.

OpenStudy (princesssleelee):

@DanJS

OpenStudy (princesssleelee):

C. The degree is 5.

OpenStudy (princesssleelee):

@mathmate

OpenStudy (princesssleelee):

@SolomonZelman

OpenStudy (prettygirl_shynice):

ill help

OpenStudy (princesssleelee):

please

OpenStudy (prettygirl_shynice):

holdon

OpenStudy (princesssleelee):

ok

OpenStudy (prettygirl_shynice):

let me know when u get the answer

OpenStudy (princesssleelee):

i dont have it

OpenStudy (prettygirl_shynice):

okay hold on

OpenStudy (prettygirl_shynice):

@pooja195

OpenStudy (princesssleelee):

@DanJS

OpenStudy (danjs):

k

OpenStudy (princesssleelee):

do you remember when you answered this question for someone else?

OpenStudy (danjs):

i think for the decending order, put the highest powers of x first

OpenStudy (princesssleelee):

descending is when the exponents decrease, from greatest to least.

OpenStudy (princesssleelee):

i know +1 is at the end, correct? because it has 0

OpenStudy (princesssleelee):

The degree is the highest exponent, therefore is the degree 5? or is '5' all of them added up?

OpenStudy (danjs):

nah i cant remember 3x2y + 4xy + 1 – 5x4y – 6y i thought the standard term order in these with x and y started with the highest power of x , then y, but if it is just the decending term order, then yeah the total of x and y powers first

OpenStudy (princesssleelee):

im so lost if you copy and paste the problem into google it pops up the one that you did... maybe if i send you the link and you look back at it you could remember?

OpenStudy (princesssleelee):

theres this one that isnt yours. http://openstudy.com/study#/updates/55cea025e4b0c5fe98060b1a

OpenStudy (princesssleelee):

http://openstudy.com/study#/updates/55c3cea7e4b0f6bb86c3bd70 this one IS yours

OpenStudy (princesssleelee):

is the first one written in descending order?

OpenStudy (danjs):

yeah i always remembered to put the highest x term first, then the y term if no x left so for the order, 3x2y + 4xy + 1 – 5x4y – 6y -5x^4*y + 3x^2*y + 4*x*y - 6y + 1

OpenStudy (danjs):

prioritize the x power, then terms in only y. then constant

OpenStudy (danjs):

the math is unchanged either way , it is just the convention people agree to list things

OpenStudy (princesssleelee):

c. is 5

OpenStudy (princesssleelee):

that is for sure, right?

OpenStudy (princesssleelee):

A.The polynomial in descending order is −5x4y+3x2y+4xy−6y+1 ???

OpenStudy (danjs):

yeah decending order of the terms, look at the x's first, put those in decending order, then next the terms in only y in decending order

OpenStudy (princesssleelee):

so C &A are right?

OpenStudy (danjs):

yaeh, the degree is the largest combined powers 5

OpenStudy (princesssleelee):

what is B?

OpenStudy (danjs):

each term is a monomial , A*x^n*y^m, the total number of them is 5

OpenStudy (princesssleelee):

I Dont get it

OpenStudy (danjs):

a constant value 1 is also 1*x^0*y^0

OpenStudy (danjs):

number of things added or subtracted together

OpenStudy (princesssleelee):

no idea lol

OpenStudy (danjs):

-5x^4*y + 3x^2*y + 4*x*y - 6y + 1 5 terms

OpenStudy (princesssleelee):

A.The polynomial in descending order is −5x4y+3x2y+4xy−6y+1 B. The polynomial by the numbers of terms is 5 C. The degree of the polynomial is 5.

OpenStudy (princesssleelee):

is that correct?

OpenStudy (danjs):

yes

OpenStudy (princesssleelee):

Let f(x) = 3x2 – x + 2 and g(x) = 5x2 – 1. Find f(g(x)). Show each step of your work. do you know this?

OpenStudy (danjs):

yeah, it is two functions compunded or chained together f(x) = 3x^2 - x + 2 means whatever you decide to make x, the function will apply that rule to it and give back the result.

OpenStudy (danjs):

f(input) = output so for the f( g(x) ) , the input of f is g(x)

OpenStudy (danjs):

f(x) = 3x2 – x + 2 and g(x) = 5x2 – 1 f ( g(x) ) = f( 5x^2 - 1)

OpenStudy (danjs):

f ( g(x) ) = f( 5x^2 - 1) = 3*( 5x^2 - 1 )^2 - (5x^2 - 1) + 2